Circular No. 6196

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

X-RAY NOVA 1994 IN SCORPIUS B. A. Harmon, M. L. McCollough, S. N. Zhang, W. S. Paciesas and

C. A. Wilson report, for the Compton Observatory/BATSE Team: “The x-ray/radio transient GRO J1655–40 is undergoing a bright outburst in hard x-rays above 20 keV. The source became detectable around July 22; its intensity exceeded 300 mCrab on July 27 and increased to 400 mCrab (±10 percent) on July 29–30. This is the brightest outburst in hard x-rays since that of February–April (IAUC 6143, 6147). The spectrum on July 29–30 extends to at least 200 keV and is well fitted by a power law with spectral index −2.4 ± 0.1.”

GRS 1915+105 S. Sazonov and R. Sunyaev, on behalf of the GRANAT/WATCH Team,

report: “The x-ray transient GRS 1915+105 was found to be bright in the observations of the GRANAT/WATCH all-sky monitor during July 19–25. The average flux in the energy band 8–20 keV was 750 ± 50 mCrab, and no significant variations in intensity were found. On July 27 the source became fainter, with flux 240 ± 120 mCrab. In the previous set of WATCH observations, during June 9–16, the source was not detected to an upper limit of 150 mCrab (2σ).

SATELLITES OF SATURN With reference to the account on IAUC 6192, R. A. Jacobson, Jet

Propulsion Laboratory, notes that the calculated angular distances from Saturn of Saturn XVIII (Pan), Saturn XV (Atlas) and Saturn XVI (Prometheus) at the time of the May 22 HST observations were −15′′.3, −18′′.9 and −19′′.3, respectively. Since these ephemerides could easily be in error at the arcsecond level, he suggests the identity of S/1995 S 1 with Pan and that of S/1995 S 2 with either Atlas or Prometheus.

COMET 67P/CHURYUMOV-GERASIMENKO CCD total R magnitudes: July 3.41 UT, 17.1 (S. M. Larson and C. W.

Hergenrother, University of Arizona Catalina Station, 1.5-m reflector; coma diameter 14′′; 24′′ tail in p.a. 247^{o}); 26.03, 17.3 (A. Galad, Astronomical and Geophysical Observatory, Comenius University, Modra, Slovakia, 0.6-m reflector); 27.05, 17.0 (Galad).

1995 August 1 Brian G. Marsden